A variable x is normally distributed with mean 22 and standard deviation 9. Round your answers to the nearest hundredth as needed.

a) Determine the z -score for x = 26 . z =

b) Determine the z -score for x = 17 . z =

c) What value of x has a z -score of 0.67 ? x =

d) What value of x has a z -score of − 0.4 ? x =

e) What value of x has a z -score of 0 ? x =
Answer #1

solution:

a

using z-score formula

z =X -\mu/ \sigma

z=26-22/9

z=0.44

b

using z-score formula

z =X -\mu/ \sigma

z=17-22/9

z=-0.56

c.

Using z-score formula  

x= z * \sigma + \mu

x= 0.67*9+22

x= 28.03

d.

Using z-score formula  

x= z * \sigma + \mu

x= -0.4 *9+22

x= 18.4

e

Using z-score formula  

x= z * \sigma + \mu

x= 0*9+22

x= 22